OpenStudy (anonymous):

A bag has 4 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick a yellow marble?

OpenStudy (sidsiddhartha):

there are total (4+3+3)=10 marbels ok

OpenStudy (sidsiddhartha):

now what is the chance of picking a red marbel from 10 marbels

OpenStudy (anonymous):

3 out of 10

OpenStudy (sidsiddhartha):

yeah

OpenStudy (anonymous):

how do I make that a different answer

OpenStudy (anonymous):

okay so a 1 out of 9 chance

OpenStudy (anonymous):

so 1 out of 10

OpenStudy (sidsiddhartha):

do u get it?

OpenStudy (anonymous):

I did, but would you mind helping me with others as well?

OpenStudy (anonymous):

A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a red marble, replacing it, and then picking a blue marble?

OpenStudy (sidsiddhartha):

so if ur gonna pick a yellow marble it will be =3/9=1/3

OpenStudy (anonymous):

my choices for an answer does not have that as a possible answer

OpenStudy (sidsiddhartha):

now for the second question u have total 15 marbel so u can choose a red in 5/15 way and then a blue in6/15 ways so required prob=(5/15)*(6/15)=6/5